OpenStudy (anonymous):

simplify. #17.) show that 5(3y-2)=15y-10 for at least 5 values of y. #18.) show that 3(8x-7)=24x-21 for at least 5 values of x. please just give me the direct anwsers for them (:

OpenStudy (anonymous):

.???

OpenStudy (riley):

Let me get a pencil to write this out. Just a min :P

OpenStudy (anonymous):

ok thanks (:

OpenStudy (anonymous):

#17.) 5(3y-2) = 5*3y-5*2 = 15y-10 for ALL values of y (eg. y=1, 2, 3, 4, 5 :-) ) #18.) same as before 3(8x-7) = 3*8x-3*7 = 24x -21 for ALL values of x

OpenStudy (riley):

Okay, for #17 what it's asking you to do is show how both of those equations equal the same thing when you plug 5 different values into them for y. The problem is asking you to "show" this. So you can go through and work each equation through with a different y value every time, like this: 5(3(1)-2) = 15(1) - 10 5(3-2) = 15-10 5(1) = 5 5 = 5 5(3(2)-2) = 15(2) - 10 5(6-2) = 30-10 5(4) = 20 20 = 20 5(3(3)-2) = 15(3)-10 5(9-2) = 45-10 5(7) = 35 35 = 35 5(3(4)-2) = 15(4)-10 5(12-2) = 60-10 5(10) = 50 50 = 50 5(3(5)-2) = 15(5)-10 5(15-2) = 75-10 5(13) = 65 65 = 65

OpenStudy (riley):

Do you get how to figure out #18 now?

OpenStudy (anonymous):

like i got the anwser for 18, but even know you did all that like its really hard for me to learn how to do these so i really dont understand and dont know how to do it ):

OpenStudy (riley):

I wish I knew a better way to explain it. Sometimes the hardest part with math is actually understanding what problems are asking for versus actually working them through. >_< You can treat #18 the same as I showed you with #17. All it's asking for you to do is to "show" how the two equations it gives you equal one another when you use a different value for y five times. So first, pick out five numbers you want to use for the y value. They can be any numbers you want. For example, in #17 I chose to use 1,2,3,4, and 5 to plug in for the y value. Then once you have your y values picked out to use. All thats left to do is plug them into the two equations and solve. If you do your math correctly you will end up with the same answer for each equation. =D